# SmoothCache: A Universal Inference Acceleration Technique for Diffusion Transformers

### v1

@@ 1.1 Related Work
Despite their successes, DiT models, like many transformer-based architectures, face challenges with computational efficiency during inference. The core operations of attention and feed-forward layers within transformers have quadratic time complexity. In the context of diffusion models, where these operations are performed repeatedly over multiple denoising steps, this quadratic dependency leads to significant computational overhead, particularly when generating high-resolution outputs or long sequences. This high inference cost poses a major challenge for deploying these models in real-world applications with limited computational resources or strict latency requirements.
@@
Efficient Diffusion Models. Diffusion models require many function evaluations to iteratively remove noise from data, and thus quickly scale in computational costs. Many previous works have focused on reducing the number of diffusion steps required to achieve high quality samples by exploring accelerated noise schedulers and fast ODE solvers [31, 17, 18, 16].Other works have focused on traditional acceleration methods for neural networks in order to reduce latency per step, such as quantization [29, 13, 7] and compression techniques [13, 27, 6]. However, the effectiveness of these techniques are limited since the loss incurred by per-step acceleration methods may become large when used in combination with optimized solvers. Addressing these challenges, a more adaptive solution could provide efficiency gains on top of timestep-reduction methods without compromising sample quality.
Diffusion Model Caching. Caching has emerged as a promising technique for accelerating DiT inference by exploiting computational redundancies in the denoising process. For U-Net based architectures, DeepCache [20] leverages cross-timestep similarities by caching and reusing up sampling feature maps, while other methods use focus on caching blocks of layers [34], cross-attention modules [15], or intermediate noise states [1]. Since these techniques rely on U-Net skip connection features not present in DiTs, further research has been conducted to apply caching to image [19, 4, 28] and video [35] DiT diffusion models. However, all the aforementioned caching techniques are designed to exploit specific principles of the candidate model or architecture, and do not generalize across different diffusion architectures and output modalities. Additionally, some techniques require large-scale training and experimental hyper-parameter searches for every sampling configuration, and are not usable out-of-box when applied to new models.
@@
SmoothCache addresses these limitations by introducing a caching approach that is simple to implement and training-free, but still makes caching decisions informed by a specific diffusion process, outperforming uniform schedules. The model-agnostic framework adapts to various model, sampling, and solver configurations with just one calibration inference pass and a single hyper-parameter $\alpha$, to achieve optimal performance.
@@ 2.2 SmoothCache
The objective of SmoothCache is to provide a training-free, model-agnostic strategy of caching and reusing layer outputs such that minimal error is introduced. However, we believe a single optimal static scheme that caches cross-timestep layer output may not exist across different model architectures, solvers, and modalities. For example, when examining the average representation error between layer outputs of consecutive time steps as shown in Fig. 2, we observe that layers generally exhibit higher differences in later time steps in the DiT-XL label-to-image model, while layers in the Open-Sora text-to-video model are more sensitive in the first and last diffusion time steps. This discovery emphasizes a need to apply the similarity principle in a generalizable technique, such that different models benefit differently based on error curves.
@@
Figure 4: SmoothCache-Eligible Layers Compute Composition of candidate models. These are computed from the MACs of the default configurations experimented on in this paper.
@@
A hyper parameter search to find all $\alpha_{i_{j}}$ has a exponential search space based on the number of layers and significantly costly. In order to simplify the caching problem, we define a single hyper parameter $\alpha>0$ to guide caching for all layers. We define $\mathcal{L}$ as the average L1 relative error of all $N$ layers of type $i$, which is selected in order to compare true representation errors between layers for all types $i$ and and positional depths $j$ in the network. Additionally, we recognize caching specific layers can introduce errors in future layers of the same type in the network. For example, if we approximate $L_{i_{j-1},t}$ with $L_{i_{j-1},t+k}$, this may introduce noise such that the calibration error $\mathcal{L}(\tilde{L}_{i_{j},t},\tilde{L}_{i_{j},t+k})$ no longer correctly approximates the true error $\mathcal{L}(L_{i_{j},t},L_{i_{j},t+k})$, which leads to poor caching decisions. In order to mitigate the cascading impact of caching layers, we group caching decisions for all layers of type $i$, such that all $j$ layers in $L_{i,t+k}$ approximate $L_{i,t}$. Thus, cached output $L_{i_{j},t+k}$ is used in place of computing $L_{i_{j},t}$ when
@@
Using cached outputs is computationally inexpensive, and allows for significant speedup of the diffusion inference process. As Eq. 4 makes no prior assumptions about the specific properties of the particular diffusion process being cached, it can be applied across multiple architectures and modalities. Additionally, because caching decisions are only dependent on calibration error, they do not change at model runtime. This ensures compatibility with existing graph compilation optimizations.
@@ 3.1 Experiment Setup
Text to Image Generation Firstly, we select the original DiT-XL-256x256 label to image diffusion model, using the original released model weights  [^2] https://github.com/facebookresearch/DiT. Since the DiT family of models is trained on the ImageNet1k dataset, we generate 50,000 256x256 images for evaluation. We test SmoothCache with the DDIM solver, and use classifier free guidance with a scale of 1.5.
@@
Evaluation Metrics We use a variety of standard metrics in each domain to demonstrate the efficiency to quality trade off. We record the Multiply-Accumulate Operations (MACs) and latency of the full diffusion process. We derive the acceleration ratio from baseline sampling latencies. To measure generation quality, we evaluate the candidate models on common metrics in the corresponding domain. For DiT-XL-256x256, we generate 50,000 images with the specified configuration and report the FID, IS, and sFID. For OpenSORA, we evaluate the performance of the model using VBench[9], and generate 946 videos based on the VBench suite prompts. We report the final scaled VBench score. Lastly, for Stable Audio Open, we use the exact same evaluation protocol described in the Stable Audio Open technical report [5] using the same evaluation code [^3] https://github.com/Stability-AI/stable-audio-metrics, reporting CLAP, FDOpenL3 and KLPaSST metrics for AudioCaps [10], MusicCaps without singing prompts [2], and Song Describer without singing prompts [22]. All speed results are measured on a single H100-80G GPU, and averaged across 50 runs.
Implementation Details As mentioned previously, we apply caching to layers that precede residual connections. In the DiT-XL-256x256 model, this includes the Self-attention and Feed-forward modules in the DiT block. In OpenSora V1.2, we cache the temporal Self-attention, Cross-attention and Feed-forward modules as well as their equivalent spatial variants (for a total of 6 types of modules). In Stable Audio Open, we cache the Self-attention, Cross-attention and Feed-forward modules. For each architecture, solver, and preset number of diffusion steps, we obtain the representation errors $L(F_{i_{j},t},F_{i_{j},t+k})$ by computing differences in layer outputs for some random samples. For all modalities, we use 10 samples for calibration. Ablations show that the choice of the number of calibration samples does not matter much, something that can be observed in Fig. 2. We also choose which samples to generate the error curves based on whether we did conditional generation during calibration or not, which we only do for OpenSora and Stable Audio Open. We fix $k\in\{1,2,3\}$ for DiT-XL-256x256 and Stable Audio Open as we determine the layer representation error to grow too large past a difference of 4 timesteps. In OpenSORA, $k$ goes up to 5 in particular due to relatively low error in the cross-attention components of the network. For DiT, we calibrate on samples generated unconditionally using the null prompt. We calibrate OpenSora on conditionally generated 480p 2s videos with randomly sampled prompts from VidProM [33]. For Stable Audio Open, we calibrate using randomly sampled prompts from the AudioCaps validation set.
@@ 3.2 Results
We present the results of SmoothCache on DiT-XL-256x256, OpenSora, and Stable Audio Open in Tables 12, and 3 respectively. In order to provide a fair comparison between acceleration and quality trade offs, we compare the default solver used in the model, such as DDIM or DPM++, with a SmoothCache implementation with specified hyper parameters. We find two configurations with identical acceleration ratios show SmoothCache to perform better on various quality metrics, indicating superior performance over the typical acceleration/quality tradeoff. For all results, we run 5 trials and report the mean and standard deviation for each metric.
Table 1: Results For DiT-XL-256x256 on using DDIM Sampling, sorted by TMACs. Note that L2C is not training free.
@@
Table 2: Results For OpenSora on Rectified Flow.
Schedule Steps VBench (%) ($\uparrow$) TMACs Latency (s)
No Cache 30 79.36 $\pm$0.19 1612.1 28.43
Ours ($\alpha$ = 0.02) 30 78.76 $\pm$0.38 1388.5 26.57
Ours ($\alpha$ = 0.03) 30 78.10 $\pm$0.51 1321.1 26.17
@@
Figure 5: SmoothCache results on DiT-XL/2-256x256 for unconditional generation with 50 DDIM sampling steps on ImageNet-1k for thresholds 0.08 and 0.18, as well as for Static Caching.
@@
Figure 7: SmoothCache Results on OpenSora for threshold 0.03 for 2s 480p videos. We show the first, middle and last frame of each video. We use the following prompts, in order from top to bottom: (1) A wave crashes against a rocky shore, sending spray high into the air. (2) Chocolate sauce is poured slowly over a stack of fluffy pancakes. (3) The bund Shanghai, pan right
@@ 3.2.1 Quantitative results
Comparing against existing literature In order to further investigate the true effectiveness of our technique, we compare SmoothCache to existing DiT caching techniques in literature developed concurrently with SmoothCache, such Fast-Forward Caching (FORA) [28] and Learning-to-Cache (L2C) [19] for Label-to-Image generation. We note that the above methods relies on specific properties of the relevant model architecture, and does not necessarily translate across different modalities, while SmoothCache is agnostic to those considerations as it models the caching scheme directly off the observed error curves. FORA would not work in OpenSora or Stable Audio Open due to the difference in the error curves as seen in Fig. 2. We see that SmoothCache outperforms FORA across different inference times at 50 sampling steps, yielding similar performance at lower inference time or better performance with the same inference time. The one exception here is L2C, which requires leveraging the full ImageNet training set to learn a policy for a given number of sampling steps. Changing the number of sampling steps requires a full retraining, and L2C has a theoretical maximum of a 2x speedup because the caching policy is only learned with skipping every other step, limitations that SmoothCache does not have.
Examining inference speedup/quality tradeoff Stable Audio Open and DiT-XL/2-256x256 provides the highest speed/quality tradeoff compared to OpenSora, which gives around a 10% speedup latency wise and around 16-22% reduction in MACs comapred to an almost 20-60% speedup for the other modalities, whose discrepancy could be attributed by the larger overhead of non-DiT components (since inference latency is measured end-to-end). We also observe that the amount of inference speed/quality tradeoff looks directly correlated to the error deviation between calibration samples which can be seen in Fig. 2, with the higher variance in error among individual samples across timesteps for OpenSora versus that for Stable Audio Open and DiT-XL/2-256x256.
@@ 3.2.2 Qualitative results
We show visual examples for all modalities. We show 256x256 images for image generation, using null-conditional prompts. For audio, we show the log-mel spectrogram in order to visualize the waveform for conditional prompt generation across the 3 datasets we measure. For video, we show the first, middle and last frame of 2 second 480p 24 fps videos at a 9:16 aspect ratio.
We see that for DiT-XL/2-256x256 outputs in Fig. 5, even though there is a drop of FID with SmoothCache applied, that there is some noticeable difference in performance quality when there is a higher threshold applied, but there exists a threshold where the quality of the model performance is visually indistinguishable from the model without caching. This shows the fine-granularity in inference speed/quality tradeoff that SmoothCache affords, while static caching as shown in FORA can only yield the lower quality SmoothCache threshold. We observe particularly that this degradation is less pronounced in Stable Audio Open in Fig. 6, where the audio waveforms look visually identical, and when listened to do not have perceptible differences. This is consistent with the qualitative results which show only a minor degradation across all measured metrics across all datasets. For OpenSora, we see more significant differences in when caching in Fig. 7, with noticeable artifacting when the higher caching threshold is applied. This is consistent with the quantitative results, and suggests this architecture/modality is more sensitive to caching than the other modalities, which makes sense due to the complex spatial/temporal modelling that other modalities explored do not have to deal with.

### v2

@@ 1.1 Related Work
Efficient Diffusion Models. Diffusion models require many function evaluations to iteratively remove noise from data, and thus quickly scale in computational costs. Many previous works have focused on reducing the number of diffusion steps required to achieve high quality samples by exploring accelerated noise schedulers and fast ODE solvers [33, 18, 19, 17]. However, not all tasks might be suitable for faster solvers with less steps, as they might rely on inversion for tasks like image-editing [25], and while work has been done to extend this to other solvers [9, 35], future solvers will face similar issues with adoption. Traditional acceleration methods such as such as quantization [31, 14, 7] and compression/distillation techniques [14, 29, 6] could improve model performance, but these techniques often require extensive retraining, or specific architectures and parameterization around a specific task in order to realize efficiency gains.
Diffusion Model Caching. Caching has emerged as a promising technique for accelerating DiT inference by exploiting computational redundancies in the denoising process. For U-Net based architectures, DeepCache [21] leverages cross-timestep similarities by caching and reusing up sampling feature maps, while other methods use focus on caching blocks of layers [38], cross-attention modules [16], or intermediate noise states [1]. However, in addition to exploiting specific properties of the U-Net architecture, these models only perform these analyses on image diffusion models, leveraging biases and assumptions in those tasks that might not necessarily generalize to other modalities. For Diffusion Transformers, further research has been conducted to apply caching to image [20, 4, 30] and video [40] DiT diffusion models. Again, these methods only study how their methods accelerate diffusion inference in their specific modalities, with very little attempt to show that these methods generalize beyond their modality of interest. FORA [30] for example might have extensive acceleration gains in image diffusion, but our preliminary investigations shows that it does not work on Audio or Video diffusion tasks. Likewise, Pyramid-Attention Broadcast [40] leverages specific qualities of video diffusion, such as its cache-able cross-attention layer in order to accelerate diffusion, on top of other GPU parallelization tricks, but Fig. 2 shows this assumption breaks in other modalities. Moreover, not all caching methods are post-training. Training-based caching methods like Learning-to-Cache [20] that require further training for specific diffusion step configurations, which is not viable in cases where access to source training data might not be available, such as in open source models trained on proprietary (non-public) data.
@@
SmoothCache attempts to address these limitations by introducing a training-free caching approach, that makes no underlying assumption about the specific task or modality at hand, outperforming uniform schedules and even matching or beating handcrafted caching techniques reported in literature. We also present multi-modal results to show that this technique does generalize across modalities, and how it adapts to various model, sampling, and solver configurations with just one calibration inference pass and a single hyper-parameter $\alpha$.
@@ 2.2 SmoothCache
The objective of SmoothCache is to provide a training-free, model-agnostic strategy of caching and reusing layer outputs such that minimal error is introduced. However, we believe a single optimal static scheme that caches cross-timestep layer output may not exist across different model architectures, solvers, and modalities. For example, we examine the average representation error between layer outputs of consecutive time steps as shown in Fig. 2. For layer output $L$ at timestep $t$ the representation error $k$ timesteps behind is defined as $E(L_{t},L_{t+k})=\frac{\|L_{t}-L_{t+k}\|_{1}}{\|L_{t}\|_{1}}$.
We observe that layers generally exhibit higher differences in later time steps in the DiT-XL label-to-image model, while layers in the Open-Sora text-to-video model are more sensitive in the first and last diffusion time steps. This discovery emphasizes a need to apply the similarity principle in a generalizable technique, such that different models benefit differently based on error curves.
@@
Figure 4: Layer Compute Composition of candidate models. These are computed from the MACs of the default model configurations without SmoothCache applied. Note that in all candidate models, SmoothCache eligible layers comprise at least 90% of compute time.
@@
A hyper parameter search to find all $\alpha_{i_{j}}$ has a exponential search space based on the number of layers and is significantly costly. In order to simplify the caching problem, we define a single hyper parameter $\alpha>0$ to guide caching for all layers. We define $\mathcal{L}$ as the average L1 relative error of all $N$ layers of type $i$, which is selected in order to compare true representation errors between layers for all types $i$ and and positional depths $j$ in the network. Additionally, we recognize caching specific layers can introduce errors in future layers of the same type in the network. For example, if we approximate $L_{i_{j-1},t}$ with $L_{i_{j-1},t+k}$, this may introduce noise such that the calibration error $\mathcal{L}(\tilde{L}_{i_{j},t},\tilde{L}_{i_{j},t+k})$ no longer correctly approximates the true error $\mathcal{L}(L_{i_{j},t},L_{i_{j},t+k})$, which leads to poor caching decisions. In order to mitigate the cascading impact of caching layers, we group caching decisions for all layers of type $i$, such that all $j$ layers in $L_{i,t+k}$ approximate $L_{i,t}$. Thus, cached output $L_{i_{j},t+k}$ is used in place of computing $L_{i_{j},t}$ when
@@
Using cached outputs is computationally inexpensive, and allows for significant speedup of the diffusion inference process. As Eq. 4 makes no prior assumptions about the specific properties of the particular diffusion process being cached, it can be applied across multiple architectures and modalities, only requiring a brief linear search to find an optimal $\alpha$ hyperparameter for each sampling configuration. Additionally, because caching decisions are only dependent on calibration error, they do not change at model runtime. This ensures compatibility with existing graph compilation optimizations.
@@ 3.1 Experiment Setup
Text to Image Generation Firstly, we select the original DiT-XL-256$\times$256 label to image diffusion model, using the original released model weights  [^2] https://github.com/facebookresearch/DiT. Since the DiT family of models is trained on the ImageNet1k dataset, we generate 50,000 256$\times$256 images for evaluation. We test SmoothCache with the DDIM solver, and use classifier free guidance with a scale of 1.5.
@@
Evaluation Metrics We use a variety of standard metrics in each domain to demonstrate the efficiency to quality trade off. We record the Multiply-Accumulate Operations (MACs) and latency of the full diffusion process. We derive the acceleration ratio from baseline sampling latencies. To measure generation quality, we evaluate the candidate models on common metrics in the corresponding domain. For DiT-XL-256$\times$256, we generate 50,000 images with the specified configuration and report the FID, IS, and sFID. For OpenSORA, we evaluate the performance of the model using VBench [10], Learned Perceptual Image Patch Similarity (LPIPS) [39], Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [37], and generate 946 videos based on the VBench suite prompts. LPIPS, PSNR and SSIM are computed relative to the non-cached videos. We report the final scaled VBench score. Lastly, for Stable Audio Open, we use the exact same evaluation protocol described in the Stable Audio Open technical report [5] using the same evaluation code [^3] https://github.com/Stability-AI/stable-audio-metrics, reporting CLAP, FDOpenL3 and KLPaSST metrics for AudioCaps [11], MusicCaps without singing prompts [2], and Song Describer without singing prompts [23]. All speed results are measured on a single H100-80G GPU, and averaged across 50 runs.
Implementation Details As mentioned previously, we apply caching to layers that precede residual connections. In the DiT-XL-256$\times$256 model, this includes the Self-attention and Feed-forward modules in the DiT block. In OpenSora V1.2, we cache the temporal Self-attention, Cross-attention and Feed-forward modules as well as their equivalent spatial variants (for a total of 6 types of modules). In Stable Audio Open, we cache the Self-attention, Cross-attention and Feed-forward modules. For each architecture, solver, and preset number of diffusion steps, we obtain the representation errors $L(F_{i_{j},t},F_{i_{j},t+k})$ by computing differences in layer outputs for some random samples. For all modalities, we use 10 samples for calibration. Ablations show that the choice of the number of calibration samples does not matter much, something that can be observed in Fig. 2. We also choose which samples to generate the error curves based on whether we did conditional generation during calibration or not, which we only do for OpenSora and Stable Audio Open. We fix $k\in\{1,2,3\}$ for DiT-XL-256$\times$256 and Stable Audio Open as we determine the layer representation error to grow too large past a difference of 4 timesteps. In OpenSORA, $k$ goes up to 5 in particular due to relatively low error in the cross-attention components of the network. For DiT, we calibrate on samples generated unconditionally using the null prompt. We calibrate OpenSora on conditionally generated 480p 2s videos with randomly sampled prompts from VidProM [36]. For Stable Audio Open, we calibrate using randomly sampled prompts from the AudioCaps validation set.
@@ 3.2 Results
We present the results of SmoothCache on DiT-XL-256$\times$256, OpenSora, and Stable Audio Open in Tables 12, and 3 respectively. In order to provide a fair comparison between acceleration and quality trade offs, we compare the default solver used in the model, such as DDIM or DPM++, with a SmoothCache implementation with specified hyper parameters. We find two configurations with identical acceleration ratios show SmoothCache to perform better on various quality metrics, indicating superior performance over the typical acceleration/quality tradeoff. For all results, we run 5 trials and report the mean and standard deviation for each metric.
Table 1: Results For DiT-XL-256$\times$256 on using DDIM Sampling, sorted by TMACs. Note that L2C is not training free.
@@
Table 2: Results For OpenSora on Rectified Flow (30 steps).
Schedule VBench (%) ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) TMACs Latency (s)
No Cache 79.36 $\pm$0.19 - - - 1612.1 28.43
Ours ($\alpha$ = 0.02) 78.76 $\pm$0.38 0.5852 $\pm$ 0.0352 11.08 $\pm$ 0.96 0.5085 $\pm$ 0.0315 1388.5 26.57
Ours ($\alpha$ = 0.03) 78.10 $\pm$0.51 0.5347 $\pm$ 0.1119 12.62 $\pm$ 2.81 0.5601 $\pm$ 0.0812 1321.1 26.17
@@
Figure 5: SmoothCache results on DiT-XL/2-256$\times$256 for unconditional generation with 50 DDIM sampling steps on ImageNet-1k for thresholds 0.08 and 0.18, as well as for Static Caching.
@@
Figure 7: SmoothCache Results on OpenSora for threshold 0.03 for 2s 480p videos. We show the first, middle and last frame of each video. We use the following prompts, in order from top to bottom: (1) Chocolate sauce is poured slowly over a stack of fluffy pancakes. (2) an astronaut floating in space with the earth in the background (3) The bund Shanghai, pan right
@@ 3.2.1 Quantitative results
Comparing against existing literature In order to further investigate the true effectiveness of our technique, we compare SmoothCache to existing DiT caching techniques in literature developed concurrently with SmoothCache, such Fast-Forward Caching (FORA) [30] and Learning-to-Cache (L2C) [20] for Label-to-Image generation. We note that the above methods relies on specific properties of the relevant model architecture, and does not necessarily translate across different modalities, while SmoothCache is agnostic to those considerations as it models the caching scheme directly off the observed error curves. FORA does not work in OpenSora or Stable Audio Open due to the difference in the error curves as seen in Fig. 2 and hence we do not report the results here. We see that SmoothCache outperforms FORA across different inference times at 50 sampling steps, yielding similar performance at lower inference time or better performance with the same inference time. The one exception here is L2C, which requires leveraging the full ImageNet training set to learn a policy for a given number of sampling steps. Changing the number of sampling steps requires a full retraining, and L2C has a theoretical maximum of a 2$\times$ speedup because the caching policy is only learned with skipping every other step, limitations that SmoothCache does not have.
Examining inference speedup/quality tradeoff Stable Audio Open and DiT-XL/2-256$\times$256 provides the highest speed/quality tradeoff compared to OpenSora, which gives around a 10% speedup latency wise and around 16-22% reduction in MACs comapred to an almost 20-60% speedup for the other modalities, whose discrepancy could be attributed by the larger overhead of non-DiT components (since inference latency is measured end-to-end). We also observe that the amount of inference speed/quality tradeoff looks directly correlated to the error deviation between calibration samples which can be seen in Fig. 2, with the higher variance in error among individual samples across timesteps for OpenSora versus that for Stable Audio Open and DiT-XL/2-256$\times$256.
@@ 3.2.2 Qualitative results
We show visual examples for all modalities. We show 256$\times$256 images for image generation, using null-conditional prompts. For audio, we show the log-mel spectrogram in order to visualize the waveform for conditional prompt generation across the 3 datasets we measure. For video, we show the first, middle and last frame of 2 second 480p 24 fps videos at a 9:16 aspect ratio.
We see that for DiT-XL/2-256$\times$256 outputs in Fig. 5, even though there is a drop of FID with SmoothCache applied, that there is some noticeable difference in performance quality when there is a higher threshold applied, but there exists a threshold where the quality of the model performance is visually indistinguishable from the model without caching. This shows the fine-granularity in inference speed/quality tradeoff that SmoothCache affords, while static caching as shown in FORA can only yield the lower quality SmoothCache threshold. We observe particularly that this degradation is less pronounced in Stable Audio Open in Fig. 6, where the audio waveforms look visually identical, and when listened to do not have perceptible differences. This is consistent with the qualitative results which show only a minor degradation across all measured metrics across all datasets. For OpenSora, we see more significant differences in when caching in Fig. 7, with noticeable artifacting when the higher caching threshold is applied. This is consistent with the quantitative results, and suggests this architecture/modality is more sensitive to caching than the other modalities, which makes sense due to the complex spatial/temporal modelling that other modalities explored do not have to deal with.